OpenStudy (anonymous):

The first five terms in a quadratic sequence are shown below. 6, 21, 46, 81, 126,... What is the nth term in the sequence? A. n2 + 5 B. 5n2 + 1 C. 6n2 + 1 D. 5n2 + 2

OpenStudy (anonymous):

If we look at B \[ n = 1 \implies 5 \cdot 1^2 + 1 = 6 \\ n = 2 \implies 5 \cdot 2^2 + 1 = 5 \cdot 4 + 1 = 21\\ n = 3 \implies 5 \cdot 3^2 + 1 = 5 \cdot 9 + 1 = 46\\ n = 4 \implies 5 \cdot 4^2 + 1 = 5 \cdot 16 + 1 = 81\\ n = 5 \implies 5 \cdot 5^2 + 1 = 5 \cdot 25 + 1 = 126\\ \]

OpenStudy (anonymous):

Thanks! it was correct.